OpenStudy (anonymous):

the cost of manufacturing clocks is given by c=16 sqrt (n+49), where c is the total cost and n is the number produced. what is c when no clocks are produced?

OpenStudy (anonymous):

when n = 0 c = 16 * 7 = 112

OpenStudy (anonymous):

n = 0 c = 16 * 7 = 112